# Decomposing The Dark Matter of Sparse Autoencoders

### v1

@@ 4 Predicting SAE Error
5 Modeling Activations
@@
We will adopt the weak linear hypothesis (Smith, 2024b), a generalization of the linear representation hypothesis which holds only that some features in language models are represented linearly. Thus we have
@@
5.1 Analyzing Error Norm Prediction
We will first analyze Eq. 2, the learned probe from ${\mathbf{x}}$ to $\lVert\texttt{SaeError}({\mathbf{x}})\rVert_{2}^{2}$. First, we claim that given a vector ${\mathbf{x}}$, if ${\mathbf{x}}$ is a sparse sum of orthogonal vectors, then there exists a perfect prediction vector ${\bm{a}}$ such that ${\bm{a}}^{T}{\mathbf{x}}\approx\left\lvert{\mathbf{x}}\right\rvert_{2}^{2}$ (in other words, the norm squared of ${\mathbf{x}}$ can be linearly predicted from ${\mathbf{x}}$). The proof of this claim is in Section B.1; the intuition is that we can set the probe vector ${\bm{a}}^{*}$ to the sum of the vectors ${\bm{y}}_{i}$ weighted by their average weight $E({\textnormal{w}}_{i})$.
When ${\mathbf{x}}$ is instead a sparse sum of non-orthogonal vectors, as it partly is in Eq. 6 amd Eq. 9, this proof is no longer true, but we now argue that a similar intuition holds. If the ${\bm{y}}_{i}$ are almost orthogonal and do not activate much at the same time, then a probe vector again equal to the sum of vectors ${\bm{y}}_{i}$ weighted by their average value $E({\textnormal{w}}_{i})$ will be a good approximate prediction. Indeed, when we try predicting $\lVert\texttt{Sae}({\mathbf{x}})\rVert_{2}^{2}$ from $\texttt{Sae}({\mathbf{x}})$ (which is a sparse sum of known almost orthogonal vectors of a similar distribution to the true SAE vectors), we find that indeed the linear probe that is learned is approximately equal to this sum (see Section B.2).
@@ 5.1 Analyzing Error Norm Prediction
5.2 Analyzing Error Vector Prediction
We will now analyze Eq. 3, the learned transformation from ${\mathbf{x}}$ to $\texttt{SaeError}({\mathbf{x}})$, with our model of SAE error from Eq. 9. We assume that $\texttt{Introduced}({\mathbf{x}})$ cannot be approximated at all as a linear function of ${\mathbf{x}}$; if this assumption is violated in practice, it will show up in our estimations as an increased amount of $\texttt{Dense}({\mathbf{x}})$.
If $\texttt{Dense}({\mathbf{x}})+\sum_{i=m}^{n}{\textnormal{w}}_{i}{\bm{y}}_{i}$ is contained in a linear subspace of ${\mathbf{x}}$ orthogonal to $\sum_{i=0}^{m}{\textnormal{w}}_{i}{\bm{y}}_{i}$, then the error of the transformation ${\bm{b}}^{*}$ exactly equals $\texttt{Introduced}({\mathbf{x}})$ (since the transformation is just exactly this orthogonal linear subspace). However, if such a linear transform does not exist, the percent of variance left unexplained by the regression will be an upper bound on the true variance explained by $\texttt{Introduced}({\mathbf{x}})$. We also note that if this test is accurate, we can use it to estimate $\texttt{Dense}({\mathbf{x}})$: the difference between the variance explained by $\texttt{Sae}({\mathbf{x}})$ and the variance explained by ${\mathbf{x}}-(\texttt{Sae}({\mathbf{x}})+{\mathbf{b}}^{*}\cdot{\mathbf{x}})$ will approach $\texttt{Dense}({\mathbf{x}})$ as $m\rightarrow\infty$.
@@ 5.2 Analyzing Error Vector Prediction
We run this synthetic setup with a Gemma Scope layer $20$ SAE (width $16k$, $L_{0}\approx 68$) in Section B.3, and find that indeed, estimated Dense(${\mathbf{x}}^{\prime}$) is highly correlated with the amount of Gaussian nosie added to ${\mathbf{x}}^{\prime}$ and Introduced(${\mathbf{x}}^{\prime}$) is highly correlated with the amount of Gaussian noise added to $\texttt{Sae}({\mathbf{x}}^{\prime})$ (see Section 5.2). However, note that because ${\mathbf{x}}^{\prime}$ noise is also slightly correlated with estimated Introduced(${\mathbf{x}}^{\prime}$), it is possible that some of the contribution to the estimated nonlinear error is from Dense(${\mathbf{x}}^{\prime}$).
@@
Thus, we again now have a potential explanation for our initial results: we can predict error vectors because they consist in large part of not yet learned linear features in an almost orthogonal subspace of ${\mathbf{x}}$, we can predict a smaller portion of larger SAE errors because the number of these linear features go down with SAE width, and the horizontal line in Fig. 1 is because $\texttt{Dense}({\mathbf{x}})$ and $\texttt{Introduced}({\mathbf{x}})$ are mostly constant. Furthermore, we can hypothesize from the correlations om Section 5.2 that the linearly predictable component of SAE error consists mostly of not yet learned features and $\texttt{Dense}({\mathbf{x}})$, while the component that is not linearly predictable consists mostly of $\texttt{Introduced}({\mathbf{x}})$. We will explore this hypothesis in Section 6.
5.3 Analyzing Per-Token Scaling Predictions

### v2

@@ 4.4 Predicting SAE Per-Token Error Norms
5 Analyzing Components of SAE Error
In previous sections, we broke down the SAE error vector into a linearly predictable component and a non-linearly predictable component. A reasonable question is whether these error subsets meaningfully differ. Thus, in this section, we run experiments on these components to determine how the linear and non-linear components of $\texttt{SaeError}({\mathbf{x}})$ differ. For convenience, given a probe ${\bm{b}}^{*}$ from Eq. 11, we write
$\texttt{LinearError}({\mathbf{x}})$ :-$\coloneq{\bm{b}}^{*}\cdot{\mathbf{x}}:$ (14)
$\texttt{NonlinearError}({\mathbf{x}})$ :-$\coloneq\texttt{SaeError}({\mathbf{x}})-\texttt{LinearError}({\mathbf{x}})=\texttt{SaeError}({\mathbf{x}})-{\bm{b}}^{*}:$ (15)
In Section B.2 we discuss a hypothesis for SAE feature activations that might explain why we observe these differences; this appendix section is not necessary for understanding the experiments in this section, and we will refer to it in the few cases where it provides additional intuition.
We run the following experiments in this section to understand how the linear and nonlinear components of SAE error differ:
  1. 1.
    In Section 5.1, we run the norm prediction test from Eq. 10 on different components of the error. We find that the norm of the nonlinear component is less predictable from activations than other components, implying that it might consist of fewer not-yet-learned features.
  2. 2.
    We train new SAEs directly on the linear and nonlinear components of error. The SAE trained on nonlinear error converges to higher reconstruction loss and produces less interpretable features.
  3. 3.
    We examine how much each component contributes to downstream model performance by intervening in the forward pass, and find that both components contribute proportionally to their size to increased cross entropy loss.
Overall, these experiments suggest that our split of SAE dark matter into these two categories is indeed a meaningful one.
Figure 6: Violin plot of norm prediction tests for all SAEs (layer $12$ Gemma 2B SAEs, layer $20$ Gemma 9B SAEs, all Gemma Scope SAEs with $L_{0}$ closest to $60$ across layers, and all Llama Scope SAEs). We plot the $R^{2}$ of a linear regression from ${\mathbf{x}}$ to to each random vector’s norm squared.
Figure 7: Auto-interpretability results on SAEs trained on the linear and nonlinear components of $\texttt{SaeError}({\mathbf{x}})$ on a width 16k, $L_{0}\approx 60$, layer $20$ Gemma Scope SAE. “Not” represents contexts that the SAE latent did not activate on, while each $Qi$ represents activating examples from decile $i$.
5.1 Applying the Norm Prediction Test:
For our first experiment, we run the norm prediction test from Eq. 10 on five different random vectors: ${\mathbf{x}}$, $\texttt{Sae}({\mathbf{x}})$, $\texttt{SaeError}({\mathbf{x}})$, $\texttt{LinearError}({\mathbf{x}})$, and $\texttt{NonlinearError}({\mathbf{x}})$. The results are shown as a violin plot for each component across all $329$ SAEs we experiment with in Section 4 (layer $12$ Gemma 2B SAEs, layer $20$ Gemma 9B SAEs, all Gemma Scope SAEs with $L_{0}$ closest to $60$ across layers, and all Llama Scope SAEs) in Fig. 7 (the $\texttt{Sae}({\mathbf{x}})$ bar can be seen as a summary of Fig. 2). There are one or two SAE with an outlier $R^{2}$ equal to or lower than 00; we omit these from the plot because they are likely due to numeric instability in the linear regression routine.
The intuition behind this test is that if a random vector consists of a sum of almost-orthogonal linear features from ${\mathbf{x}}$, then its norm should be linearly predictable from ${\mathbf{x}}$ (see Appendix B for more on this intuition).
Firstly, we note that $\left\lVert{\mathbf{x}}\right\rVert_{2}^{2}$ can almost be perfectly predicted from ${\mathbf{x}}$. This is reassuring news for the linear representation hypothesis, as it implies that ${\mathbf{x}}$ may be well modeled as the sum of many one-dimensional features, at least from the perspective of this test. We also find that $\texttt{NonlinearError}({\mathbf{x}})$ (and $\texttt{SaeError}({\mathbf{x}})$, which consists partly of $\texttt{NonlinearError}({\mathbf{x}})$) has a notably lower score on this test than $\texttt{LinearError}({\mathbf{x}})$. This result suggests that $\texttt{NonlinearError}({\mathbf{x}})$ does not consist of a sparse sum of linear features from ${\mathbf{x}}$.
5.2 Training SAEs on $\texttt{SaeError}({\mathbf{x}})$ Components:
Another empirical test we run is training an SAE on $\texttt{NonlinearError}({\mathbf{x}})$ and $\texttt{LinearError}({\mathbf{x}})$. We choose a fixed Gemma 9B Gemma Scope layer $20$ SAE with $16k$ latents and $L_{0}\approx 60$ to generate $\texttt{SaeError}({\mathbf{x}})$ from. This SAE has nonlinear and linear components of the error approximately equal in norm and $R^{2}$ of the total $\texttt{SaeError}({\mathbf{x}})$ they explain, so it presents a fair comparison. We train SAEs to convergence (about 100M tokens) on each of these components of error and find that the SAE trained on $\texttt{NonlinearError}({\mathbf{x}})$ converges to a fraction of variance unexplained an absolute $5$ percent higher than the SAE trained on the linear component of SAE error ($\approx 0.59$ and $\approx 0.54$ respectively).
We additionally examine the interpretability of the learned SAE latents using automated interpretability (this technique was first proposed by Bills et al. (2023) for interpreting neurons, and first applied to SAEs by Cunningham et al. (2023)). Specifically, we use the implementation introduced by Juang et al. (2024), where a language model (we use Llama 3.1 70b (AI@Meta, 2024)) is given top activating examples to generate an explanation, and then must use only that explanation to predict if the feature fires on a test context. Our results in Fig. 7 show that the SAE trained on linear error produces latents that are about an absolute $5\%$ more interpretable across all activation firing deciles (we average results across $1000$ random features for both SAEs, where for each feature use $7$ examples in each of the $10$ feature activation deciles as well as $50$ negative examples, and show $95\%$ confidence intervals).
Figure 8: Results of intervening in the forward pass and replacing ${\mathbf{x}}$ with $\texttt{Sae}({\mathbf{x}})+\texttt{NonlinearError}({\mathbf{x}})$ and $\texttt{Sae}({\mathbf{x}})+\texttt{LinearError}({\mathbf{x}})$ during the forward pass on all layer $20$ Gemma 9B SAEs with $L_{0}\approx 60$ or width of 16k. Reported in percent of cross entropy loss recovered with respect to the difference between the same intervention with $\texttt{Sae}({\mathbf{x}})$ and with the normal model forward pass.
5.3 Downstream Cross Entropy Loss of $\texttt{SaeError}({\mathbf{x}})$ Components:
A common metric used to test SAEs is the percent of cross entropy loss recovered when the SAE reconstruction is inserted into the model in place of the original activation versus an ablation baseline (see e.g. Bloom (2024)). We modify this test to specifically examine the different components of $\texttt{SaeError}({\mathbf{x}})$: we compare the percent of the cross entropy loss recovered when replacing ${\mathbf{x}}$ with $\texttt{Sae}({\mathbf{x}})$ plus either $\texttt{LinearError}({\mathbf{x}})$ or $\texttt{NonlinearError}({\mathbf{x}})$ to the baseline of inserting just $\texttt{Sae}({\mathbf{x}})$ in place of ${\mathbf{x}}$ on all layer $20$ Gemma 9B SAEs with $L_{0}\approx 60$ or width of 16k. To estimate how much each component “should” recover, we use two metrics: the average norm of the component relative to the total norm of $\texttt{Sae}({\mathbf{x}})$ and the percent of the variance that the component recovers between $\texttt{Sae}({\mathbf{x}})$ and ${\mathbf{x}}$. The results, shown in Fig. 8, show that for the most part these metrics are reasonable predictions for both types of error. That is, both $\texttt{NonlinearError}({\mathbf{x}})$ and $\texttt{LinearError}({\mathbf{x}})$ proportionally contribute to the SAE’s increase in downstream cross entropy loss, with possibly a slightly higher contribution than expected for $\texttt{LinearError}({\mathbf{x}})$.
6 Reducing $\texttt{NonlinearError}({\mathbf{x}})$
In past sections we identify $\texttt{NonlinearError}({\mathbf{x}})$ and $\texttt{LinearError}({\mathbf{x}})$ and argue that $\texttt{NonlinearError}({\mathbf{x}})$ likely does not consist of not-yet-learned linear features. This may be a problem for scaling SAEs farther: if a part of SAE error persistently consists of nonlinear transformations of linear features, these may never be able to be learned by SAEs. Thus, in this section, we investigate two approaches for reducing $\texttt{NonlinearError}({\mathbf{x}})$:
  1. 1.
    Using a more powerful encoder at inference time. This approach has limited success, reducing total FVU by 3-5% but leaving $\texttt{FVU}_{\texttt{nonlinear}}$ almost unchanged
  2. 2.
    Attempting to predict $\texttt{NonlinearError}({\mathbf{x}})$ from SAE reconstructions of adjacent model components - this approach is more successful; on the SAE we test on, $\texttt{Sae}({\mathbf{x}})$ predicts 50% of the nonlinear error, and previous model components predict between 4% and 14% of nonlinear error, although only between 6% and 10% of the overall SAE error.
6.1 Using a More Powerful Encoder
Figure 9: Investigations towards reducing nonlinear SAE error.
Our first approach for reducing nonlinear error is to try improving the encoder. We use a recent approach suggested by Smith (2024a): applying a greedy inference time optimization (ITO) algorithm called gradient pursuit to a frozen learned SAE decoder matrix. We implement and run ITO on all layer $20$ Gemma Scope $9b$ SAEs closest to $L_{0}\approx 60$. For each example ${\mathbf{x}}$ with reconstruction $\texttt{Sae}({\mathbf{x}})$, we use the gradient pursuit implementation with an $L_{0}$ exactly equal to the $L_{0}$ of ${\mathbf{x}}$ in the original $\texttt{Sae}({\mathbf{x}})$.
Using these new reconstructions of ${\mathbf{x}}$, we repeat Eq. 11 and do a linear transformation from ${\mathbf{x}}$ to the inference time optimized reconstructions. We then regenerate the same scaling plot as Fig. 1 and show this figure in Fig. 9(a). Our first finding is that pursuit indeed decreases the total FVU of $\texttt{Sae}({\mathbf{x}})$ by $3$ to $5\%$; as Smith (2024a) only showed an improvement on a small $1$ layer model, to the best of our knowledge we are the first to show this result on state of the art SAEs. Our most interesting finding, however, is that the $\texttt{FVU}_{\texttt{nonlinear}}$ stays almost constant when compared to the original SAE scaling in Fig. 1. We hypothesize that this might happen because ITO reduces “easy” linear errors like feature shrinkage. In Fig. 1, we plot the additional reduction in $\texttt{FVU}_{\texttt{nonlinear}}$ as the contribution of encoder error; because $\texttt{FVU}_{\texttt{nonlinear}}$ stays almost constant, this section is very narrow.
6.2 Predicting SAE Errors Between SAEs:
For the Gemma 2 architecture at the locations the SAEs are trained on, each residual activation can be decomposed in terms of prior components:
We focus on $layer=20$ and Gemma 2 9B, and thus use the layer $19$ attention and MLP Gemma Scope SAEs. For all components we use the width $16k$ SAEs with $L_{0}\approx 60$. In Fig. 9(b), we plot the $R^{2}$ of a regression from the SAE output corresponding to each of these right hand side components to each of the different components of an SAE trained on $\texttt{Resid}_{\texttt{layer}}$ ($\texttt{SaeError}({\mathbf{x}})$, $\texttt{LinearError}({\mathbf{x}})$, and $\texttt{NonlinearError}({\mathbf{x}})$).
We find that we can explain a small amount (up to $\approx 10\%$) of total $\texttt{SaeError}({\mathbf{x}})$ using previous components, which may be immediately useful for circuit analysis. We also can explain large parts of $\texttt{NonlinearError}({\mathbf{x}})$ with prior components as well. While $\texttt{Sae}({\mathbf{x}})$ explains $50\%$ of the variance in the nonlinear error, this may be somewhat misleading, as the nonlinear error is partially a function of $\texttt{Sae}({\mathbf{x}})$:
These results mean that we might be able to explain some of the SAE Error using a circuits level view, but that overall there are still large parts of each error component unexplained.
7 Conclusion
The fact that SAE error can be predicted and analyzed at all is surprising. Thus, our findings are intriguing evidence that SAE errors, and not just SAE reconstructions, are worthy of analysis, and we hope that it inspires further work in decomposing SAE error.
Concretely, we believe that our study of error has already discovered a number of promising directions for future SAE research. The discovery that SAE errors can be significantly predicted from model activations hints that some subspaces may resist being learned and points toward potential architectural innovations such as incorporating low-rank dense side channels. The predictable relationship between small and large SAE errors may also streamline experimentation with novel SAE architectures by allowing researchers to efficiently forecast scaling behavior through small-scale trials. Additionally, our demonstration that prior SAE outputs can explain part of SAE error has immediate practical implications for circuit analysis, which currently relies on large noise terms that complicate circuit interpretation.
At a higher level, the presence of constant nonlinear error at a fixed sparsity as we scale implies that scaling SAEs may not be the only (or best) way to explain more of model behavior. Future work might explore alternative penalties besides sparsity or new ways to learn better dictionaries. Ultimately, we believe that there is still room to make SAEs better, not just bigger.
Acknowledgments
Appendix A Extra Error Prediction Experiments
A.1 Note on Feature Shrinkage
Earlier SAE variants were prone to feature shrinkage: the observation that $\texttt{Sae}({\mathbf{x}})$ systematically undershot ${\mathbf{x}}$. Current state of the art SAE variants (e.g. JumpReLU SAEs and TopK SAEs, which we examine in this work), are less vulnerable to this problem, although we still find that Gemma Scope reconstructions have about a $10$% smaller norm than ${\mathbf{x}}$. One potential concern is that the ${\mathbf{b}}^{*}$ in Eq. 11 that we learn is merely predicting this shrinkage. If this was the case, then the cosine similarity of the linear error prediction $({\bm{b}}^{*})^{T}\cdot{\mathbf{x}}$ with ${\mathbf{x}}$ would be close to $1$; however, in practice we find that it is around $0.5$, so ${\bm{b}}^{*}$ is indeed doing more than predicting shrinkage.
A.2 Additional Error Prediction Plots
In Fig. 10(b) and Fig. 11(b), we show the accuracy of SAE error norm predictions, SAE error vector predictions, and $\texttt{FVU}_{\texttt{nonlinear}}$ for layers 5 and 19 of Gemma 2 2B and layers 9 and 31 of Gemma 2 9B. As described in the main text, we find broadly similar results for SAE error norm and error vector prediction at these layers, but find that $\texttt{FVU}_{\texttt{nonlinear}}$ decreases some as we increase SAE width at a fixed $L_{0}$, although this result is uncertain because of the sparsity of Gemma Scope SAEs at these layers.
A.3 Norm Prediction Baselines
In Fig. 14, we run a linear regression from different model variables to $\lVert\texttt{SaeError}({\mathbf{x}})\rVert$ across layers on Gemma Scope 9B for SAEs of size 131k. We find that is is not “easy” to predict SAE error norm, especially at later layers; the token identity, SAE L0, activation norm, and model loss all do significantly worse than using the full activation. It is interesting to note that at the first few layers, token identity does better at predicting SAE error than a probe of the activations; this is perhaps not surprising, since recent results from e.g. Lad et al. (2024) show that very early layers primarily operate on a per token level.
A.4 Breaking Apart Error Per Token
Appendix B Modeling Activations
@@
In this section, we seek to explain why we see a difference between linear and nonlinear error in the main body of the paper. We will adopt the weak linear hypothesis (Smith, 2024b), a generalization of the linear representation hypothesis which holds only that some features in language models are represented linearly. Thus we have
@@
Before we proceed with our analysis, we note that there is a rich sparse coding literature studying dictionary learning in the setting with mixed sparse and dense signals. For example, in a classic work Candès et al. (2011) propose Robust Principal Component Analysis for decomposing data matrices into sparse and dense components, although this is not directly applicable to our setting of trying to learn an autoencoder. More recently, Tasissa et al. propose a dictionary learning technique for data exactly modeled as in Eq. 17. In our work, we focus on studying existing SAEs and hypothesizing why they fail, so these works are not immediately applicable, but we are excited to see future work that applies these combined sparse-dense autoencoders to language model activations..
@@
B.1 Analyzing Error Norm Prediction
We will first analyze Eq. 10, the learned probe from ${\mathbf{x}}$ to $\lVert\texttt{SaeError}({\mathbf{x}})\rVert_{2}^{2}$. First, we claim that given a vector ${\mathbf{x}}$, if ${\mathbf{x}}$ is a sparse sum of orthogonal vectors, then by basic linear algebra there exists a perfect prediction vector ${\bm{a}}$ such that ${\bm{a}}^{T}{\mathbf{x}}\approx\left\lvert{\mathbf{x}}\right\rvert_{2}^{2}$ (in other words, the norm squared of ${\mathbf{x}}$ can be linearly predicted from ${\mathbf{x}}$). The proof of this claim is in Section C.1; the intuition is that we can set the probe vector ${\bm{a}}^{*}$ to the sum of the vectors ${\bm{y}}_{i}$ weighted by their average weight $E({\textnormal{w}}_{i})$.
When ${\mathbf{x}}$ is instead a sparse sum of non-orthogonal vectors, as it partly is in Eq. 17 and Eq. 20, this proof is no longer true, but we now argue that a similar intuition holds. If the ${\bm{y}}_{i}$ are almost orthogonal (formally $coherence<\epsilon$, see (Foucart & Rauhut, 2013)) and do not activate much at the same time, then a probe vector again equal to the sum of vectors ${\bm{y}}_{i}$ weighted by their average value $E({\textnormal{w}}_{i})$ will be a good approximate prediction. Indeed, when we try predicting $\lVert\texttt{Sae}({\mathbf{x}})\rVert_{2}^{2}$ from $\texttt{Sae}({\mathbf{x}})$ (which is a sparse sum of known almost orthogonal vectors of a similar distribution to the true SAE vectors), we find that indeed the linear probe that is learned is approximately equal to this sum (see Section C.2).
@@ B.1 Analyzing Error Norm Prediction
B.2 Analyzing Error Vector Prediction
We will now analyze Eq. 11, the learned transformation from ${\mathbf{x}}$ to $\texttt{SaeError}({\mathbf{x}})$, with our model of SAE error from Eq. 20. We assume that $\texttt{Introduced}({\mathbf{x}})$ cannot be approximated at all as a linear function of ${\mathbf{x}}$. This is a reasonable assumption since the SAE is a nonlinear function of ${\mathbf{x}}$, but if indeed some of $\texttt{Introduced}({\mathbf{x}})$ can be approximated in this way, then we will underestimate the amount of $\texttt{Introduced}({\mathbf{x}})$ and therefore overestimate the amount of $\texttt{Dense}({\mathbf{x}})$.
If $\texttt{Dense}({\mathbf{x}})+\sum_{i=m}^{n}{\textnormal{w}}_{i}{\bm{y}}_{i}$ is contained in a linear subspace of ${\mathbf{x}}$ orthogonal to $\sum_{i=0}^{m}{\textnormal{w}}_{i}{\bm{y}}_{i}$, then $\texttt{Dense}({\mathbf{x}})$ is part of the linearly explainable error, and the error of the transformation ${\bm{b}}^{*}$ exactly equals $\texttt{Introduced}({\mathbf{x}})$ (since the transformation is just exactly this orthogonal linear subspace). However, if such an orthogonal linear subspace does not exist, the optimal linear transform will only be able to reconstruct part of $\texttt{Dense}({\mathbf{x}})+\sum_{i=m}^{n}{\textnormal{w}}_{i}{\bm{y}}_{i}$, and the percent of variance left unexplained by the regression will be an upper bound on the true variance explained by $\texttt{Introduced}({\mathbf{x}})$. We also note that if this linear transform indeed recovers $\texttt{Dense}({\mathbf{x}})$ and absent features but fails to recover $\texttt{Introduced}({\mathbf{x}})$, we can use it to estimate $\texttt{Dense}({\mathbf{x}})$: the difference between the variance explained by $\texttt{Sae}({\mathbf{x}})$ and the variance explained by ${\mathbf{x}}-(\texttt{Sae}({\mathbf{x}})+{\mathbf{b}}^{*}\cdot{\mathbf{x}})$ will approach $\texttt{Dense}({\mathbf{x}})$ as $m\rightarrow\infty$.
@@ B.2 Analyzing Error Vector Prediction
We run this synthetic setup with a Gemma Scope layer $20$ SAE (width $16k$, $L_{0}\approx 68$) in Section C.3, and find that indeed, estimated Dense(${\mathbf{x}}^{\prime}$) is highly correlated with the amount of Gaussian noise added to ${\mathbf{x}}^{\prime}$ and Introduced(${\mathbf{x}}^{\prime}$) is highly correlated with the amount of Gaussian noise added to $\texttt{Sae}({\mathbf{x}}^{\prime})$ (see Section B.2). However, note that because ${\mathbf{x}}^{\prime}$ noise is also slightly correlated with estimated Introduced(${\mathbf{x}}^{\prime}$), it is possible that some of the contribution to the estimated nonlinear error is from Dense(${\mathbf{x}}^{\prime}$).
@@
Thus, we again now have a potential explanation for our initial results: we can predict error vectors because they consist in large part of not yet learned linear features in an almost orthogonal subspace of ${\mathbf{x}}$, we can predict a smaller portion of larger SAE errors because the number of these linear features go down with SAE width, and the horizontal line in Fig. 1 is because $\texttt{Dense}({\mathbf{x}})$ and $\texttt{Introduced}({\mathbf{x}})$ are mostly constant. Furthermore, we can hypothesize from the correlations om Section B.2 that the linearly predictable component of SAE error consists mostly of not yet learned features and $\texttt{Dense}({\mathbf{x}})$, while the component that is not linearly predictable consists mostly of $\texttt{Introduced}({\mathbf{x}})$. We will explore this hypothesis in Section 5.
B.3 Analyzing Per-Token Scaling Predictions